# The Converse Madelung Question Schrödinger Equation from Minimal Axioms

### v1

@@ Abstract
We define and address the converse Madelung question: not whether Fisher information can reproduce quantum mechanics, but whether it is necessary. We adopt minimal, physically motivated axioms on hydrodynamic variables: locality, probability conservation, Euclidean invariance with global $\mathrm{U}(1)$ phase symmetry, reversibility, and convex regularity. Within the ensuing class of first-order local Hamiltonian field theories, the Poisson bracket is uniquely fixed to the canonical bracket on $(\rho,S)$ under the Dubrovin-Novikov hypotheses for local first-order hydrodynamic brackets with probability conservation. Under a pointwise, gauge-covariant complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$, among convex, rotationally invariant, first-derivative local functionals of $\rho$, whose Euler-Lagrange contribution yields a reversible completion that becomes exactly projectively linear is the Fisher functional. With $\hbar^{2}=2m\alpha$ the dynamics reduce to the linear Schrödinger equation. In many-body systems, exact projective linearity with a single local complex structure across all sectors forces $\alpha_{i}=\hbar^{2}/(2m_{i})$ componentwise, thereby fixing a single Planck constant. Galilean covariance appears via the Bargmann central extension in this framework, with the usual superselection implications. Comparison with the Doebner-Goldin family identifies the reversible corner at zero diffusion as the linear Schrödinger case in our variables. We supply operational falsifiers via residual diagnostics for the continuity and Hamilton-Jacobi equations and report numerical minima at the Fisher scale that are invariant under Galilean boosts. These results are consistent with viewing quantum mechanics, in this setting, as a reversible fixed point of Fisher-regularised information hydrodynamics. A code archive accompanies the work for direct numerical verification and reproducibility, including a superposition stress-test showing that, in our tested families and to numerical precision under grid refinement, the Fisher regulariser preserves exact projective linearity within our axioms.
@@ 1  Introduction
We ask whether Fisher curvature is not only sufficient but necessary within a strictly local, first order, reversible Hamiltonian class on $(\rho,S)$ endowed with Euclidean invariance, global $\mathrm{U}(1)$ phase symmetry, and convex regularity in $\rho$ alone. Formally: classify all admissible first derivative convex regularisers $f(\rho,\nabla\rho)$ and compatible Poisson brackets on $(\rho,S)$ for which there exists a pointwise, derivative-free, gauge-covariant complexifier that renders the time evolution exactly projectively linear on rays. The claim established here is a uniqueness within this admissible class, not a statement about higher-derivative, weakly nonlocal, mixed $(\rho,S)$, or open-system extensions.
Our scope is strictly local and first order in spatial derivatives on $(\rho,S)$; nonlocal terms, higher-derivative regularisers, and open-system couplings lie outside the present analysis. Reversibility together with parity excludes any explicit $S$-dependence in the convex regulariser and rules out dissipative couplings between $\rho$ and $S$. As an operational falsifier, after the complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ we require exact preservation of $\psi\mapsto a\,\psi_{1}+b\,\psi_{2}$ under time stepping and we quantify any deviation by a residual norm; this test complements, but does not replace, the analytic uniqueness proof. Within this setting we show that the Dubrovin-Novikov bracket classification reduces the hydrodynamic bracket to the canonical form on $(\rho,S)$, and that the only admissible first-derivative convex regulariser of $\rho$ that yields a reversible completion compatible with exact projective linearity is the Fisher functional. For contrast with the Doebner-Goldin diffusion families [4, 5], note that the diffusive sector lies outside our reversible Hamiltonian cone; in the zero-diffusion corner one recovers the linear Schrödinger equation in our variables, while nonlinear gauge-equivalent representatives fail the exact projective linearity stress-test except in the Fisher case.
This gap between sufficiency and necessity motivates the central question of this work. Does an alternative, non-Fisher regulariser exist that also satisfies reversibility, locality, and exact projective linearity within the stated class? We refer to this investigation into necessity as the converse Madelung question.
@@ 2  Minimal Axioms
We restrict attention to first-order local hydrodynamic Poisson structures of Dubrovin-Novikov (DN) type [13, 14] acting on scalar doublets $u=(\rho,S)$. This ensures the reversible Poisson operator is the flat representative, ruling out derivative-dependent coefficients and zeroth-order cores under the stated symmetries. The Poisson-operator coefficients depend only on the local fields and not on their derivatives; by Euclidean invariance together with Jacobi they are constant, ensuring both reversibility and well-posedness. Throughout we take $\Omega=\mathbb{T}^{d}$ (periodic) or $\mathbb{R}^{d}$ with standard decay so that $\rho,|\nabla S|\!\to\!0$ as $|x|\!\to\!\infty$.
@@
We adopt the local canonical Poisson bracket
@@
with $\{\rho,\rho\}=0=\{S,S\}$. All time evolutions are $\dot{F}=\{F,\mathcal{H}\}$ for a real Hamiltonian $\mathcal{H}[\rho,S]$.
@@
belongs to $H^{-1}$ even at nodal sets [3, 9]. Boundary terms vanish under the conditions listed in Appendix A; admissible classes are periodic, Dirichlet with $S$ constant on $\partial\Omega$, or Neumann with vanishing normal derivatives.
@@
Dubrovin-Novikov locality. By DN type we mean first-order local hydrodynamic operators acting on the scalar doublet $u=(\rho,S)$, whose coefficients depend on the fields but not their derivatives. Translation and rotation invariance, the presence of the conserved phase generator $\mathcal{C}=\!\int\rho\,dx$, and Jacobi restrict the admissible class to a flat, constant-coefficient representative that is Poisson-isomorphic to the canonical bracket written above. We therefore work, without loss within our axioms, with the canonical form.
@@
The restriction to first-order Dubrovin-Novikov brackets ensures that the flow of $\rho$ is local and divergence-free. Higher-order Hamiltonian operators, such as third-order form in the KdV hierarchy, either introduce additional dimensional parameters or violate probability conservation by producing higher-derivative fluxes that are not expressible as $\nabla\!\cdot j$. The first-order case is therefore the minimal setting in which locality and the Jacobi identity can coexist for a probability field.
Appendix C shows that, once Euclidean covariance and reversibility are imposed, attempts to include derivative dependence beyond first order in a local scalar operator on $(\rho,S)$ either violate the Jacobi identity or introduce extra Casimirs. Within these constraints, the Dubrovin-Novikov first-order class appears to be the appropriate maximal setting for our constraints.
@@
Axiom II requires $\partial_{t}\rho=-\,\nabla\!\cdot(\rho\,\nabla S/m)$ for arbitrary states, hence $2\,a(\rho)=\rho/m$ and
@@
Reversibility means Hamiltonian flow with an antisymmetric bracket and a real Hamiltonian, (see Appendix O). The evolution is generated by an antisymmetric bilinear bracket satisfying Jacobi:
@@
Algebraic reversibility refers to Hamiltonian flow generated by an antisymmetric bracket with a real $\mathcal{H}$; physical time-reversal invariance here is the symmetry $(t,S)\mapsto(-t,-S)$ of the equations. For the canonical bracket and $\mathcal{H}[\rho,|\nabla S|^{2},\rho\text{-only terms}]$ these coincide. Any diffusive addition, e.g. $\varepsilon\,\Delta\rho$ in the continuity equation with $\varepsilon>0$, breaks both and yields non-negative Shannon entropy production $dS_{\mathrm{Sh}}/dt=\varepsilon\!\int|\nabla\rho|^{2}/\rho\,dx\geq 0$.
@@
$Q_{f}(\rho)\equiv-\,\frac{1}{2}\,\frac{\delta F}{\delta\rho}.$
Within this class, Proposition 2.3 (proved in Appendix D) shows that the unique choice compatible with exact projective linearity after a local complexifier is $f(\rho)=\kappa/\rho$, i.e.
$F[\rho]=4\kappa\!\int_{\Omega}|\nabla\sqrt{\rho}|^{2}dx,\qquad Q_{\kappa}(\rho)=-\,\kappa\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}.$
Any other $f(\rho)$ produces a residual nonlinear term in the Hamilton-Jacobi equation that cannot be removed by any local, gauge-preserving change of variables within our admissible class.
@@
For $\rho(x)=\exp(-x^{2}/\sigma^{2})$ and $F[\rho]=\int\alpha\,\frac{|\nabla\rho|^{2}}{\rho}\,dx$, write $\rho=R^{2}$ so $F=4\alpha\int|\nabla R|^{2}dx$. Then $\delta F/\delta\rho=-\,\alpha\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}$ in $H^{-1}$, which produces the Laplacian quotient used in Section 2. An identical calculation holds for a compactly supported bump, with nodes excluded by a standard mask window as detailed in Appendix E.
@@
Under Axioms I-V and the existence of the conserved phase generator $\mathcal{C}=\!\int\rho\,dx$ with $\{S,\mathcal{C}\}=-1$, any first-order DN-type Poisson operator on $(\rho,S)$ is Poisson-isomorphic to the canonical bracket $\{\rho(x),S(y)\}=\delta(x-y)$.
@@
Let $F[\rho]=\int f(\rho)\,|\nabla\rho|^{2}dx$ be as in Axiom VI. If there exists a local gauge-covariant complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ that maps the Hamiltonian flow generated by $\mathcal{H}=\int\frac{\rho|\nabla S|^{2}}{2m}+V\rho+Q_{f}(\rho)\,dx$ to a linear unitary evolution on $L^{2}$ for all admissible states, then $f(\rho)=\kappa/\rho$ and the resulting evolution is the linear Schrödinger equation with $\hbar^{2}=2m\kappa$.
@@ 2.1 Interdependence
  • Without VI: any other $f(\rho)$ gives a nonlinear Schrödinger equation.
@@
Axioms A1-A6 $\Rightarrow$ canonical bracket on $(\rho,S)$ $\Rightarrow$ Fisher curvature by convexity and symmetry $\Rightarrow$ local complexifier rigidity $\Rightarrow$ linear Schrödinger dynamics with fixed scale, all within the stated class of local first-derivative Hamiltonian theories.
A summary of explicit one-line counterexamples illustrating logical independence is provided in Appendix B; each axiom can fail while the others hold.
@@ 3  Bracket Classification
Lemma 3.3 (Jacobi for $\delta$ only brackets).
For $\{F,G\}_{a_{0}}=\int a_{0}(\rho)\big(F_{\rho}G_{S}-F_{S}G_{\rho}\big)\,dx$ the Jacobi identity holds if and only if $a_{0}^{\prime}(\rho)=0$, i.e. $a_{0}$ is constant.
@@
Compute the Schouten bracket $P^{i\ell}\partial_{\ell}P^{jk}+\text{cyclic}=0$ for $P^{\rho S}=a_{0}(\rho)=-P^{S\rho}$ and other entries zero. The only nonvanishing derivative is $\partial_{\rho}P^{\rho S}=a_{0}^{\prime}(\rho)$, and the $(\rho,S,\rho)$ component yields $a_{0}\,a_{0}^{\prime}=0$ pointwise, hence $a_{0}^{\prime}=0$. ∎
Thus $a_{0}$ is a constant that rescales $S$; set $a_{0}=1$.
The constant $a_{0}$ merely rescales $S$; we set $a_{0}=1$ without loss of generality. Appendix C gives the explicit trilinear Jacobi calculation indicating this result, including the mixed permutation check.
@@
with $(g^{ij},b^{ij}_{\;k})$ satisfying flatness and compatibility conditions equivalent to Jacobi [13, 14]. Translation and rotation invariance, together with the existence of the phase generator $\mathcal{C}$ and Lemma 3.3, imply that within our axioms the bracket is Poisson isomorphic to a flat constant representative, namely the canonical bracket $\{\rho(x),S(y)\}=\delta(x-y)$. Details are given in Appendix C.
@@
Equivalently, within the Dubrovin-Novikov class the bracket reduces (up to a constant rescaling of $S$) to this constant-coefficient normal form; any $\rho$-dependent prefactor $a_{0}(\rho)$ in $\{\rho,S\}=a_{0}(\rho)$ violates the Jacobi identity, hence the canonical bracket (3.4) is the unique local representative in this setting.
The proof follows from the gauge generator normal form, the $\mathrm{U}(1)$ restriction, Lemma 3.3, and the DN reduction above.
@@
This global topology affects boundary conditions but leaves the local canonical bracket unchanged.
The Jacobi identity therefore fixes the Poisson structure up to an overall constant multiplier. Allowing $a_{0}^{\prime}(\rho)\neq 0$ introduces curvature in the phase-space measure and destroys the Lie-Poisson property.

### v2

@@ Abstract
We define and address the converse Madelung question: not whether Fisher information can reproduce quantum mechanics, but whether it is necessary. We adopt minimal, physically motivated axioms on hydrodynamic variables: locality, probability conservation, Euclidean invariance with global $\mathrm{U}(1)$ phase symmetry, reversibility, and convex regularity. Within the ensuing explicitly restricted class of first-order local Hamiltonian field theories, the Poisson bracket is uniquely fixed to the canonical bracket on $(\rho,S)$ under the Dubrovin-Novikov hypotheses for local first-order hydrodynamic brackets with probability conservation. Under a pointwise, gauge-covariant complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$, among convex, rotationally invariant, first-derivative local functionals of $\rho$, the only one whose Euler-Lagrange contribution yields a reversible completion that becomes exactly projectively linear is the Fisher functional. With $\hbar^{2}=2m\alpha$ the dynamics reduce to the linear Schrödinger equation. In many-body systems, exact projective linearity with a single local complex structure across all sectors forces $\alpha_{i}=\hbar^{2}/(2m_{i})$ componentwise, thereby fixing a single Planck constant. Galilean covariance appears via the Bargmann central extension in this framework, with the usual superselection implications. Comparison with the Doebner-Goldin family identifies the reversible corner at zero diffusion as the linear Schrödinger case in our variables. We supply operational falsifiers via residual diagnostics for the continuity and Hamilton-Jacobi equations and report numerical minima at the Fisher scale that are invariant under Galilean boosts. These results are consistent with viewing quantum mechanics, in this setting, as a reversible fixed point of Fisher-regularised information hydrodynamics. A code archive accompanies the work for direct numerical verification and reproducibility, including a superposition stress-test showing that, in our tested families and to numerical precision under grid refinement, the Fisher regulariser preserves exact projective linearity within our axioms.
@@ 1  Introduction
We ask whether Fisher curvature is not only sufficient but necessary within a strictly local, first order, reversible Hamiltonian class on $(\rho,S)$ endowed with Euclidean invariance, global $\mathrm{U}(1)$ phase symmetry, and convex regularity in $\rho$ alone. Formally: classify all admissible first derivative convex regularisers $f(\rho,\nabla\rho)$ and compatible Poisson brackets on $(\rho,S)$ for which there exists a pointwise, derivative-free, gauge-covariant complexifier that renders the time evolution exactly projectively linear on rays. In less formal terms, we ask whether any other local first-derivative curvature on $\rho$ can support a reversible hydrodynamic completion that still looks exactly linear at the $\psi$ level after a single local complexification.
The claim established here is a uniqueness within this admissible class, not a statement about higher-derivative, weakly nonlocal, mixed $(\rho,S)$, or open-system extensions.
1.1 Scope
Our scope is strictly local and first order in spatial derivatives on $(\rho,S)$; nonlocal terms, higher-derivative regularisers, and open-system couplings lie outside the present analysis. Reversibility together with parity excludes any explicit $S$-dependence in the convex regulariser and rules out dissipative couplings between $\rho$ and $S$. As an operational falsifier, after the complexifier $\psi=\sqrt{\rho}\,e^{iS/\hbar}$ we require exact preservation of $\psi\mapsto a\,\psi_{1}+b\,\psi_{2}$ under time stepping and we quantify any deviation by a residual norm; this test complements, but does not replace, the analytic uniqueness proof.
Within this setting we show that the Dubrovin-Novikov bracket classification reduces the hydrodynamic bracket to the canonical form on $(\rho,S)$, and that the only admissible first-derivative convex regulariser of $\rho$ that yields a reversible completion compatible with exact projective linearity is the Fisher functional.
For contrast with the Doebner-Goldin diffusion families [4, 5], note that the diffusive sector lies outside our reversible Hamiltonian cone; in the zero-diffusion corner one recovers the linear Schrödinger equation in our variables, while nonlinear gauge-equivalent representatives fail the exact projective linearity stress-test except in the Fisher case.
This gap between sufficiency and necessity motivates the central question of this work. Does an alternative, non-Fisher regulariser exist that also satisfies reversibility, locality, and exact projective linearity within the stated class? We refer to this investigation into necessity as the converse Madelung question. Within the axioms stated below the answer will be negative: once locality, reversibility and projective linearity are imposed simultaneously, the Fisher functional is the only surviving curvature in this class.
@@ 2  Minimal Axioms
This section defines the local reversible Hamiltonian class that we work in throughout. We collect our axioms and treat them as a fixed setting. All later uniqueness and necessity statements are made inside this class, rather than for arbitrary models on wave functions or densities.
Sections 3 - 8 then form a single chain: the axioms first fix the bracket, then constrain the Hamiltonian and curvature, and finally isolate the unique local complexification that yields linear Schrödinger dynamics.
@@
We restrict attention to first-order local hydrodynamic Poisson structures of Dubrovin-Novikov (DN) type [13, 14] acting on scalar doublets $u=(\rho,S)$. This ensures the reversible Poisson operator is represented within a flat DN class, ruling out derivative-dependent coefficients and zeroth-order cores under the stated symmetries. The Poisson-operator coefficients depend only on the local fields and not on their derivatives; Euclidean invariance together with the DN hypotheses restricts them to a flat, constant-coefficient representative (classified in Appendix C), ensuring both reversibility and well-posedness.
Throughout we take $\Omega=\mathbb{T}^{d}$ (periodic) or $\mathbb{R}^{d}$ with standard decay so that $F[\rho]<\infty$ and $Q_{\alpha}\in H^{-1}(\Omega)$ even in the presence of nodal sets.
@@
Anticipating the Dubrovin-Novikov classification in Proposition 2.2, we work within our axioms in the canonical local representative
@@
belongs to $H^{-1}$ even at nodal sets [3, 9]. All Hamilton-Jacobi equations and residual diagnostics in this paper use this sign convention for $Q_{\alpha}$. Boundary terms vanish under the conditions listed in Appendix A; admissible classes are periodic, Dirichlet with $S$ constant on $\partial\Omega$, or Neumann with vanishing normal derivatives. As shown in Proposition 2.2 below, under Axioms I-V any admissible Dubrovin-Novikov bracket on $(\rho,S)$ is Poisson-isomorphic to this canonical form, so no generality is lost by adopting it here.
@@
Dubrovin-Novikov locality. As already stated above, by DN type we mean first-order local hydrodynamic operators acting on the scalar doublet $u=(\rho,S)$, whose coefficients depend on the fields but not their derivatives. Translation and rotation invariance, together with the presence of the conserved phase generator $\mathcal{C}=\!\int\rho\,dx$ and the Jacobi identity in the DN class, restrict the admissible brackets to a flat, constant-coefficient representative that is Poisson-isomorphic to the canonical bracket written above (see Appendix C for the classification). We therefore work, without loss within our axioms, with the canonical form.
@@
The restriction to first-order Dubrovin-Novikov brackets ensures that the flow of $\rho$ is local and divergence-free. Higher-order Hamiltonian operators, such as the third-order form in the KdV hierarchy, either introduce additional dimensional parameters or violate probability conservation by producing higher-derivative fluxes that are not expressible as $\nabla\!\cdot j$. The first-order case is therefore the minimal setting in which locality and the Jacobi identity can coexist for a probability field.
Appendix C shows that, once Euclidean covariance and reversibility are imposed, attempts to include derivative dependence beyond first order in a local scalar operator on $(\rho,S)$ either violate the Jacobi identity or introduce extra Casimirs. Within these constraints, the Dubrovin-Novikov first-order class appears to be the appropriate maximal setting for our constraints. We make no claim to classify Dubrovin-Novikov brackets in full generality, only those compatible with this scalar hydrodynamic, locality and symmetry ansatz.
@@
Within our class, Axiom II fixes the probability current to be $j=\rho\,\nabla S/m$, so probability conservation requires $\partial_{t}\rho=-\,\nabla\!\cdot(\rho\,\nabla S/m)$ for arbitrary states; hence $2\,a(\rho)=\rho/m$ and
@@
Reversibility means Hamiltonian flow with an antisymmetric bracket and a real Hamiltonian (see Appendix O). The evolution is generated by an antisymmetric bilinear bracket satisfying Jacobi:
@@
Algebraic reversibility refers to Hamiltonian flow generated by an antisymmetric bracket with a real $\mathcal{H}$; physical time-reversal invariance here is the symmetry $(t,S)\mapsto(-t,-S)$ of the equations. For the canonical bracket and $\mathcal{H}[\rho,|\nabla S|^{2},\rho\text{-only terms}]$ these coincide. Any diffusive addition, e.g. $\varepsilon\,\Delta\rho$ in the continuity equation with $\varepsilon>0$, breaks both and yields non-negative Shannon entropy production for $S_{\mathrm{Sh}}[\rho]:=-\int_{\Omega}\rho\ln\rho\,dx$, namely $dS_{\mathrm{Sh}}/dt=\varepsilon\!\int_{\Omega}|\nabla\rho|^{2}/\rho\,dx\geq 0$.
@@
$Q_{f}(\rho)\equiv\frac{1}{2}\,\frac{\delta F}{\delta\rho}.$
Within this class we will show, via Proposition 2.3 (proved in Appendix D), that the unique choice compatible with exact projective linearity after a local complexifier is $f(\rho)=\kappa/\rho$, i.e.
$F[\rho]=4\kappa\!\int_{\Omega}|\nabla\sqrt{\rho}|^{2}dx,\qquad Q_{\kappa}(\rho)=-\,2\,\kappa\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}.$
Any other $f(\rho)$ produces a residual nonlinear term in the Hamilton-Jacobi equation that cannot be removed by any local, gauge-preserving change of variables within our admissible class. Within the minimal convex regularity class adopted here, this is a genuine no-go statement: there is no alternative local curvature that preserves all axioms and still yields an exactly linear projective evolution.
@@
For $\rho(x)=\exp(-x^{2}/\sigma^{2})$ and $F[\rho]=\int\alpha\,\frac{|\nabla\rho|^{2}}{\rho}\,dx$, write $\rho=R^{2}$ so $F=4\alpha\int|\nabla R|^{2}dx$. Then $\delta F/\delta\rho=-\,4\alpha\,\frac{\Delta\sqrt{\rho}}{\sqrt{\rho}}$ in $H^{-1}$, which produces the Laplacian quotient used in Section 2. An identical calculation holds for a compactly supported bump, with nodes excluded by a standard mask window as detailed in Appendix E.
@@
Under Axioms I-V and the existence of the conserved phase generator
$\mathcal{C}=\int\rho\,dx,\qquad\{S,\mathcal{C}\}=-1,$
any first-order DN-type Poisson operator on $(\rho,S)$ is Poisson-isomorphic to the canonical local bracket
$\{\rho(x),S(y)\}=\delta(x-y).$
@@
Let $F[\rho]=\int f(\rho)\,|\nabla\rho|^{2}\,dx$ be as in Axiom VI. If there exists a local gauge-covariant complexifier
$\psi=\sqrt{\rho}\,e^{iS/\hbar}$
that maps the Hamiltonian flow generated by
$\mathcal{H}=\int\left[\frac{\rho|\nabla S|^{2}}{2m}+V\rho+\rho\,Q_{f}(\rho)\right]dx$
to a linear unitary evolution on $L^{2}$ for all admissible states, then $f(\rho)$ must be of Fisher form
$f(\rho)=\frac{C}{\rho},\qquad C>0,$
and the resulting evolution is the linear Schrödinger equation. For the Fisher choice $f(\rho)=C/\rho$ in Axiom VI this regulariser can equivalently be written as $F[\rho]=\alpha\int_{\Omega}|\nabla\sqrt{\rho}|^{2}dx$, so that the contribution $\rho\,Q_{f}(\rho)$ coincides, up to boundary terms, and in the variational sense, with the curvature term in the Hamiltonian density (4.1) used in Section 2.
@@ 2.1 Interdependence
  • Without VI: the convex regulariser $F[\rho]=\int f(\rho)\,|\nabla\rho|^{2}\,dx$ is no longer constrained to the Fisher form; by Proposition 2.3 any non-Fisher choice $f(\rho)$ that still admits a local complexifier produces a nonlinear Schrödinger equation within our class.
A counterexample within scope must therefore take one of two forms: a concrete error in the derivations under the axioms, or an explicit model satisfying all axioms while reproducing our diagnostics yet producing a non-Schrödinger reversible flow; models outside of axiomatic scope do not refute the classification but instead redefine the problem.
@@
Axioms I-VI $\Rightarrow$ canonical bracket on $(\rho,S)$ $\Rightarrow$ Fisher curvature by convexity and symmetry $\Rightarrow$ local complexifier rigidity $\Rightarrow$ linear Schrödinger dynamics with fixed scale, all within the stated class of local first-derivative Hamiltonian theories.
@@ 3  Bracket Classification
Lemma 3.3 (Gauge-normalised $\delta$-only brackets).
For $\{F,G\}_{a_{0}}=\int a_{0}(\rho)\big(F_{\rho}G_{S}-F_{S}G_{\rho}\big)\,dx$ with $\{\rho,\rho\}=0=\{S,S\}$, the Jacobi identity is automatically satisfied for any smooth $a_{0}(\rho)$. Axiom II and the requirement that $\mathcal{C}=\!\int\rho\,dx$ generate constant shifts of $S$ then force $a_{0}$ to be a nonzero constant, which can be absorbed into a rescaling of $S$ so that $a_{0}=1$.
@@
For a two-component local bracket with $\{\rho,\rho\}=0=\{S,S\}$ the only potentially nonzero component of the Poisson tensor is $P^{\rho S}(\rho)=a_{0}(\rho)=-P^{S\rho}$. In two dimensions the Schouten bracket $[P,P]$ vanishes identically for any such antisymmetric tensor, so the Jacobiator is identically zero and imposes no constraint on $a_{0}$. Axiom II requires $\{S,\mathcal{C}\}=-1$ with $\mathcal{C}=\!\int\rho\,dx$, which forces $a_{0}$ to be a nonzero constant; this constant can be absorbed into a redefinition of $S$, so we may set $a_{0}=1$. ∎
The constant $a_{0}$ merely rescales $S$; we set $a_{0}=1$ without loss of generality. Appendix C gives an explicit trilinear Jacobi calculation confirming that no further constraint arises from Jacobi beyond those already implied by Axiom II.
@@
with $(g^{ij},b^{ij}_{\;k})$ satisfying flatness and compatibility conditions equivalent to Jacobi [13, 14].
Translation and rotation invariance, together with the existence of the phase generator $\mathcal{C}$ and Lemma 3.3, imply that within our axioms the bracket is Poisson-isomorphic to a flat constant representative, namely the canonical bracket $\{\rho(x),S(y)\}=\delta(x-y)$. Details are given in Appendix C.
@@
Equivalently, within the Dubrovin-Novikov class the bracket reduces (up to a constant rescaling of $S$) to this constant-coefficient normal form; any $\rho$-dependent prefactor $a_{0}(\rho)$ in $\{\rho,S\}=a_{0}(\rho)$ is incompatible with the phase-generator property of Axiom II and probability conservation, hence the canonical bracket (3.4) is the unique local representative in this setting.
The proof follows from the gauge generator normal form, the $\mathrm{U}(1)$ restriction, Lemma 3.3, and the DN reduction above. In words, global phase invariance fixes the pairing of $\rho$ and $S$, isotropy removes derivative-weighted coefficients, and the Jacobi identity leaves only a constant prefactor that can be absorbed into $S$.